My daughter is having trouble with the following problem, and I can't remember how to do it, let alone explain it... Please help!
(x-5)/6 = (x-11)/2

Found 2 solutions by Nate, checkley75:
Answer by Nate(3500)

(x - 5)/6 = (x - 11)/2 mult. each side by the lowest number each one goes into
6(x - 5)/6 = 6(x - 11)/2
x - 5 = 3(x - 11)
x - 5 = 3x - 33
28 = 2x
14 = x

Answer by checkley75(3666)

(x-5)/6=(x-11)/2 now cross multiply
6(x-11)=2(x-5)
6x-66=2x-10
6x-2x=66-10
4x=56
x=56/4
x=14 answer.